OpenStudy (anonymous):

find the solution set of the equation t-√3t+6=-2

OpenStudy (anonymous):

\[t-\sqrt{3t+6}=-2\]

OpenStudy (tkhunny):

#1 Consider the Domain: \(3t + 6 \ge 0\) #2 Isolate the radical on one side.